# A Statistical Framework of Watermarks for Large Language Models: Pivot, Detection Efficiency and Optimal Rules

### v1

@@ Abstract
Since ChatGPT was introduced in November 2022, embedding (nearly) unnoticeable statistical signals into text generated by large language models (LLMs), also known as watermarking, has been used as a principled approach to provable detection of LLM-generated text from its human-written counterpart. In this paper, we introduce a general and flexible framework for reasoning about the statistical efficiency of watermarks and designing powerful detection rules. Inspired by the hypothesis testing formulation of watermark detection, our framework starts by selecting a pivotal statistic of the text and a secret key—provided by the LLM to the verifier—to enable controlling the false positive rate (the error of mistakenly detecting human-written text as LLM-generated). Next, this framework allows one to evaluate the power of watermark detection rules by obtaining a closed-form expression of the asymptotic false negative rate (the error of incorrectly classifying LLM-generated text as human-written). Our framework further reduces the problem of determining the optimal detection rule to solving a minimax optimization program. We apply this framework to two representative watermarks—one of which has been internally implemented at OpenAI—and obtain several findings that can be instrumental in guiding the practice of implementing watermarks. In particular, we derive optimal detection rules for these watermarks under our framework. These theoretically derived detection rules are demonstrated to be competitive and sometimes enjoy a higher power than existing detection approaches through numerical experiments.
@@ 1 Introduction
  3. Data pollution. Another potential concern is that the internet will soon consist of more LLM-generated text than human-written text. If LLM-generated text is indiscriminately mixed with human-written text for training, it becomes difficult to create high-quality data for developing next-generation LLMs [48, 54, 17].
An initial effort is to leverage specific patterns of LLM-generated text to distinguish it from human-generated text [25, 71, 42]. However, this approach has become increasingly ineffective as models such as ChatGPT-4, Claude 3, Gemini 1.5 Pro, and many others have reached a level that makes it significantly difficult, if not impossible, to distinguish their generated text from human-written text [65]. A more viable approach is to watermark text by embedding a signal into the LLM-generated text in a manner that allows the watermark to be provably detected [15, 1]. The property of provable detection is crucial because it allows the verifier to identify LLM-generated text for malicious purposes without relying on assumptions about the text, which may not always hold. Moreover, a reasonable watermarking scheme should be approximately unbiased, meaning it does not significantly distort the meaning or style of the original text.
@@
This reality has made it imperative for researchers to develop watermarking methods for LLM-generated text. Within a year since 2023, numerous watermarks have been proposed [34, 22, 36, 29, 67, 75, 74, 39, 24]. A common feature of these methods is leveraging the probabilistic nature of LLMs. In essence, these methods incorporate pseudorandomness into the text generation process of LLMs, and the coupling between the LLM-generated text and the pseudorandomness serves as a signal that can be used for detecting the watermark. The constructed signal becomes pronounced for detection only when the pseudorandom numbers are known, making it practically difficult for one to remove the watermark without access to the pseudorandom numbers [15].
@@
After generating text in the form of a token sequence, ${w}_{1}{w}_{2}\cdots{w}_{t-1}$, the (unwatermarked) LLM generates the next token ${w}_{t}$ according to a multinomial distribution $\bm{P}_{t}:=(P_{t,w})_{w\in{{\mathcal{W}}}}$ on the vocabulary ${{\mathcal{W}}}$, satisfying [^2] As a convention, throughout this paper three subscripts appear: $\text{letter}_{t},\text{letter}_{{w}}$, and $\text{letter}_{t,{w}}$, which correspond to step $t$, token ${w}$, and token ${w}$ at step $t$, respectively.
$\sum_{w\in{{\mathcal{W}}}}P_{t,w}=1.$
We call $\bm{P}_{t}$ the next-token prediction (NTP) distribution at step $t$, and it depends on all prior generated tokens, the user-supplied prompt, as well as system prompts that are hidden from users [62, 49, 11].
@@
With the generated text now indistinguishable from that of the unwatermarked LLM, it seems at first glance hopeless to obtain provable detectability for watermarked text [52, 72]. Interestingly, detection can be made possible by carefully designing the decoder ${\mathcal{S}}$ to impose a coupling relationship between the token and pseudorandom variable, even without knowing the NTP predictions [15]. [^4] We cannot use the probabilities $\bm{P}_{t}$’s as the verifier in general does not have access to the LLM that generates the LLM. Moreover, the prompt for generating the text is unavailable to the verifier, hence the verifier cannot obtain $\bm{P}_{t}$ even having access to the LLM. In the following, we elaborate on this point by considering perhaps the simplest example of a watermark that achieves both unbiasedness and provable detectability.
A baby watermark.
@@
which measures the correlation between the tokens and pseudorandom variables. When the watermark is present, this statistic tends to be larger in distribution than when the watermark is absent. One can conclude that a watermark is detected if this statistic is above a certain threshold. Despite being intuitive, however, this statistic is ad hoc, and one cannot rule out the possibility of a better detection rule.
To better understand practical watermarking schemes, we turn to real-world LLMs where the size of the token vocabulary is large. As a recap, at the core of an unbiased watermark is a sampling method for multinomial distributions. Perhaps the two most common sampling methods are the Gumbel-max trick [40, 31] and the inverse transform [19]. Interestingly, these two sampling methods correspond precisely to two important and representative watermarks [36, 1].
Gumbel-max watermark [1].
Let $\zeta=(U_{w})_{{w}\in{{\mathcal{W}}}}$ consist of $|{{\mathcal{W}}}|$ i.i.d. copies of $U(0,1)$. A version of the Gumbel-max trick [26, 46, 40, 31] states that
$\arg\max_{{w}\in{{\mathcal{W}}}}\frac{\log U_{w}}{P_{w}}$
follows the NTP distribution $\bm{P}\equiv(P_{{w}})_{{w}\in{{\mathcal{W}}}}$. Recognizing this fact, Scott Aaronson proposed the following decoder [1]:
@@
Inverse Transform Watermark [36].
It is well-known that any univariate distribution can be sampled by applying the inverse cumulative distribution function (CDF) to $U(0,1)$. Given an NTP distribution $\bm{P}$ and $\pi$ that maps all tokens in ${{\mathcal{W}}}$ to a permutation of $1,2,\ldots,|{{\mathcal{W}}}|$, consider the multinomial distribution with probability mass $P_{\pi^{-1}(i)}$ at $i$ for $1\leq i\leq|{{\mathcal{W}}}|$. The CDF of this distribution takes the form
$F(x;\pi)=\sum_{{w}^{\prime}\in{{\mathcal{W}}}}P_{{w}^{\prime}}\cdot{\mathbf{1}}_{\{\pi({w}^{\prime})\leq x\}}.$
Taking as input $U\sim U(0,1)$, the generalized inverse of this CDF is defined as
$F^{-1}(U;\pi)=\min\bigg{\{}i:\sum_{{w}^{\prime}\in{{\mathcal{W}}}}P_{{w}^{\prime}}\cdot{\mathbf{1}}_{\{\pi({{w}^{\prime}})\leq i\}}\geq U\bigg{\}},$
which by construction follows the multinomial distribution $\bm{P}$ after applying the permutation $\pi$. Making use of this fact, [36] proposed the inverse transform watermark with the following decoder:
@@
where the pseudorandom variable $\zeta:=(\pi,U)$ and the permutation $\pi$ is uniformly at random. By definition, this watermark is unbiased. To detect the watermark, [36] proposed detection rules that, roughly speaking, examine the absolute difference between $U_{t}$ and the rank of the generated token, $\pi_{t}({w}_{t})$, and sum the differences across the token sequence (see details in Section 4). When the text is watermarked, the difference turns out to be small because the involved two random variables are highly dependent. Similar to the Gumbel-max watermark, the detection rules proposed by [36] could potentially benefit from a more principled derivation to enhance their performance.
@@ 1.1 This Paper
From a statistical viewpoint, a watermark scheme’s performance at the detection phase is evaluated by two competing criteria. The first is how likely it would mistakenly detect human-written text as LLM-generated text, and the second is the probability that it would mistakenly classify LLM-generated text as human-written counterpart. This perspective relates watermarks to hypothesis testing, which formally calls the two aforementioned errors Type I error and Type II error, respectively. In this regard, the effectiveness of the Gumbel-max watermark and inverse transform watermark, even including the baby watermark, is not clear yet, though [22, 47] compared their efficiency through empirical experiments. For example, even though they come with detection rules that seem intuitive, it is not clear if they are statistically optimal in the sense of having the optimal trade-off between Type I and Type II errors. If they are not optimal, it is of interest to find a better detection rule.
@@ 1.2 Related Work
The most closely related work to ours is the Gumbel-max watermark [1] and the inverse transform watermark [36]. In [75, 22], the authors introduced unbiased watermarks that are robust to probability perturbations or multi-bit processing. Other unbiased watermarks in this fast-growing line of research include [15, 67, 29]. In addition, a popular example is the red-green list watermark, which splits the vocabulary into red-green lists based on hash values of previous n-grams and slightly increases the probability of green tokens embedding the watermark [33, 34, 74, 39, 12]. In the detection phase, a high frequency of green tokens suggests the text is LLM-generated. This type of watermark is biased since the NTP distributions have been altered, thereby leading to a performance degradation of the LLM.
In contrast, there is much less work on the theoretical front. One exception is [30], which conceptualized watermark detection as a problem of composite dependence testing, with the aim of understanding the corresponding statistical limit, such as the minimax Type II error and the most powerful test achieving it. A significant difference between [30] and our work is that the former assumes the NTP distributions remain unchanged from predicting the first token to the last token. This is a very restrictive assumption from a practical perspective, as we observe from experiments that the NTP distributions indeed vary substantially.
@@ 2.1 Working Hypotheses
Working Hypothesis 2.1 is grounded purely in cryptographic considerations. In cryptography, there are well-established approaches to efficiently constructing and computing the pseudorandom number as a function of text and the secret key [8, 45, 56, 53, 32]. The pseudorandom number is very sensitive to the key, making it computationally indistinguishable from its truly random counterpart without knowledge of the key. Specifically, although $\zeta_{t}$ is completely determined by the prior text ${w}_{1:(t-1)}$ (and the secret key), a run of the hash function ${\mathcal{A}}$ could effectively introduce fresh randomness, thereby making $\zeta_{t}$ statistically independent of ${w}_{1:(t-1)}$. Note that, during detection, the communication cost is merely to pass the secret key to the verifier as the prior tokens are publicly available.
@@
The two working hypotheses have been adopted by the literature, albeit not as explicit as our treatment. For example, [67, 75] assumed conditions on the pseudorandom hash function so that the first working hypothesis is valid. [^5] In particular, they assume ${\mathcal{A}}({w}_{1:n},{\mathtt{Key}})$ i.i.d. for any possible ${w}_{1:n}$. Other works directly impose distributional assumptions on some summary statistics to effectively satisfy the two working hypotheses [33, 34, 74, 22]. As a departure from these works, [36] considered a hash function of the form $\zeta_{t}={\mathcal{A}}(t,{\mathtt{Key}})$. Consequently, the independence between the pseudorandom number and prior tokens follows by construction. However, it is worthwhile mentioning that this form of the hash function hampers computational efficiency in detection due to the necessity of iterating through the entire text sequence multiple times.
@@ 2.2 Pivotal Statistics
That is, we reject that the text is written by a human if $T_{h}$ is above $\gamma_{n,\alpha}$. The threshold $\gamma_{n,\alpha}$ is chosen to ensure significance level at $\alpha$: ${\mathbb{P}}_{H_{0}}(T_{h}(Y_{1:n})=1)=\alpha$. As the null distribution of $h\circ Y$ is known, an estimator of $\gamma_{n,\alpha}$ when the text is sufficiently long is
@@ 2.3 Class-Dependent Efficiency
Once the Type I error is controlled, we seek to evaluate the Type II error and use it as a measure to ascertain which choice of score function is more desired than others. If the NTP distributions $\bm{P}_{t}$ were known and remained the same with respect to varying $t$, the optimal score function would simply be given by the log-likelihood ratio, according to the Neyman–Pearson lemma. However, this is not the case. We cope with this challenge by assuming that the NTP distributions belong to a distribution class, denoted as ${\mathcal{P}}$. The flexibility of using a distribution class lies in that one can choose a small class when much is known about the distributional properties of the LLM, and choose a large class if little is known.
@@
We call $R_{{\mathcal{P}}}(h)$ the ${\mathcal{P}}$-dependent efficiency rate of $h$, or simply the ${\mathcal{P}}$-efficiency rate.
@@
Under a certain regularity condition, (12) is tight in the sense that there exists $\bm{P}^{\star}$ in ${\mathcal{P}}$ such that, if $\bm{P}_{t}=\bm{P}^{\star}$ for all $t$, then
for any positive $\varepsilon$ and sufficiently large $n$. The regularity condition is satisfied by both the Gumbel-max and inverse transform watermarks and is detailed in the remark following the proof of Theorem 2.1 in Appendix A.
@@
This efficiency rate presents a significant departure from the Bahadur efficiency [6], though closely related. The Bahadur efficiency quantifies the rate at which the significance level ($p$-value) of a test statistic approaches zero with increasing sample size. It emphasizes the control of Type I errors for simple hypothesis testing with i.i.d. data. In contrast, our metric $R_{{\mathcal{P}}}(h)$ centers on quantifying the least decline of Type II errors to zero when the underlying NTP distribution $\bm{P}$ is picked up from the distribution class ${\mathcal{P}}$. In short, it focuses on composite hypothesis testing where the observed data is not identically distributed.
Optimality via minimax optimization.
@@
$\sup_{h}R_{{\mathcal{P}}}(h)=-\inf_{h}\inf_{\theta\geq 0}\sup_{\bm{P}\in{\mathcal{P}}}\left(\theta\operatorname{\mathbb{E}}_{0}h(Y)+\log\phi_{\bm{P},h}(\theta)\right)=-\inf_{h,\theta\geq 0}\sup_{\bm{P}\in{\mathcal{P}}}\left(\operatorname{\mathbb{E}}_{0}\theta h(Y)+\log\operatorname{\mathbb{E}}_{1,\bm{P}}\mathrm{e}^{-\theta h(Y)}\right).$
@@
Choice of distribution classes.
@@
Assume that at least $\gamma$-fraction of $\bm{P}_{1},\ldots,\bm{P}_{n}$ are in ${\mathcal{P}}_{1}$ with the rest being in ${\mathcal{P}}_{2}$, [^9] We do not need to know which are in the $\gamma$ fraction. where $10<\gamma<$ 1 and ${\mathcal{P}}_{1}\subset{\mathcal{P}}_{2}$. Then, for any $h$, we have
$\limsup\limits_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)^{1/n}\leq\mathrm{e}^{-\gamma\cdot R_{{\mathcal{P}}_{1}}(h)+(1-\gamma)\cdot R_{{\mathcal{P}}_{2}}(h)}.$
When there is no prior at all about the LLMs, the associated class is the $(|{{\mathcal{W}}}|-1)$-dimensional simplex, which contains all possible NTP distributions and we denote by ${\mathrm{Simp}}({{\mathcal{W}}})$. This distribution class includes singular distributions—a distribution $\bm{P}$ such that $P_{{w}}=1$ for some token ${w}$—for which the pivotal statistic has the same distribution under the null and alternative in (9). Formally, this gives $R_{{\mathrm{Simp}}({{\mathcal{W}}})}(h)=0$, that is, the class-dependent efficiency of ${\mathrm{Simp}}({{\mathcal{W}}})$ is zero. Taking ${\mathcal{P}}_{1}={\mathcal{P}}\subset{\mathcal{P}}_{2}={\mathrm{Simp}}({{\mathcal{W}}})$, Proposition 2.1 shows
$\lim\sup\limits_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)^{1/n}\leq\mathrm{e}^{-\gamma R_{{\mathcal{P}}}(h)}.$
This result implies that when only a fraction of NTP distributions are known to some extent, the Type II error is still governed by $R_{{\mathcal{P}}}(h)$ up to a multiplicative constant $\gamma$. The comparison between choices of the score functions $h$ can still be reduced to comparing the value of $R_{{\mathcal{P}}}(h)$.
@@
the $\Delta$-regular distribution class. Accordingly, an NTP distribution $\bm{P}$ is called $\Delta$-regular if $\bm{P}\in{\mathcal{P}}_{\Delta}$. This excludes singular distributions. Interestingly, $\Delta$-regularity is closely related to the Shannon entropy. For any $\Delta$-regular NTP distribution $\bm{P}$, its Shannon entropy satisfies
${\mathrm{Ent}}(\bm{P})=\sum P_{{w}}\log\frac{1}{P_{w}}\geq\sum P_{{w}}(1-P_{w})\geq\sum P_{{w}}\cdot\Delta=\Delta,$
where the first inequality follows because $\log\frac{1}{1-x}\geq x$ for $x<1$. This shows that $\Delta$-regularity imposes a lower bound on how much information the NTP distribution offers. In practice, most NTP distributions are $\Delta$-regular for a proper value of $\Delta$. See Figure 1 for an empirical investigation where 88.65% of the NTP distributions are 0.0001-regular, 81.36% are 0.001-regular, and 70.28% are 0.01-regular.

### v2

@@ Abstract
Since ChatGPT was introduced in November 2022, embedding (nearly) unnoticeable statistical signals into text generated by large language models (LLMs), also known as watermarking, has been used as a principled approach to provable detection of LLM-generated text from its human-written counterpart. In this paper, we introduce a general and flexible framework for reasoning about the statistical efficiency of watermarks and designing powerful detection rules. Inspired by the hypothesis testing formulation of watermark detection, our framework starts by selecting a pivotal statistic of the text and a secret key—provided by the LLM to the verifier—to control the false positive rate (the error of mistakenly detecting human-written text as LLM-generated). Next, this framework allows one to evaluate the power of watermark detection rules by obtaining a closed-form expression of the asymptotic false negative rate (the error of incorrectly classifying LLM-generated text as human-written). Our framework further reduces the problem of determining the optimal detection rule to solving a minimax optimization program. We apply this framework to two representative watermarks—one of which has been internally implemented at OpenAI—and obtain several findings that can be instrumental in guiding the practice of implementing watermarks. In particular, we derive optimal detection rules for these watermarks under our framework. These theoretically derived detection rules are demonstrated to be competitive and sometimes enjoy a higher power than existing detection approaches through numerical experiments.
@@ 1 Introduction
  3. Data pollution. The internet will soon consist of more LLM-generated text than human-written text. If LLM-generated text is indiscriminately mixed with human-written text for training, it becomes difficult to create high-quality data for developing next-generation LLMs [50, 56, 17].
An initial effort to mitigate these problems has been to leverage specific patterns of LLM-generated text to distinguish it from human-generated text [25, 74, 44]. However, this approach has become increasingly ineffective as models such as ChatGPT-4, Claude 3, Gemini 1.5 Pro, and many others have reached a level that makes it significantly difficult, if not impossible, to distinguish their generated text from human-written text [68].
A more viable approach is to watermark text by embedding a signal into the LLM-generated text in a manner that allows the watermark to be provably detected [15, 1]. The property of provable detection is crucial because it allows the verifier to identify LLM-generated text for malicious purposes without relying on assumptions about the text, which may not always hold. Moreover, a reasonable watermarking scheme should be approximately unbiased, meaning it does not significantly distort the meaning or style of the original text.
@@
This reality has made it imperative for researchers to develop watermarking methods for LLM-generated text. Within a year since 2023, numerous watermarks have been proposed [36, 22, 38, 30, 70, 78, 77, 41, 24]. A common feature of these methods is leveraging the probabilistic nature of LLMs. In essence, these methods incorporate pseudorandomness into the text-generation process of LLMs, and the coupling between the LLM-generated text and the pseudorandomness serves as a signal that can be used for detecting the watermark. The constructed signal becomes pronounced for detection only when the pseudorandom numbers are known, making it practically difficult for one to remove the watermark without access to the pseudorandom numbers [15].
@@
[…]
@@
With the generated text now indistinguishable from that of the unwatermarked LLM, it seems at first glance hopeless to obtain provable detectability for watermarked text [54, 75]. Interestingly, detection can be made possible by carefully designing the decoder ${\mathcal{S}}$ to impose a coupling relationship between the token and pseudorandom variable, even without knowing the NTP predictions [15]. [^4] We cannot use the probabilities $\bm{P}_{t}$’s as the verifier in general does not have access to the LLM that generates the tokens. Moreover, the prompt for generating the text is unavailable to the verifier, hence the verifier cannot obtain $\bm{P}_{t}$ even having access to the LLM. In the following, we elaborate on this point by considering perhaps the simplest example of a watermark that achieves both unbiasedness and provable detectability.
A baby watermark
@@
which measures the correlation between the tokens and pseudorandom variables. When the watermark is present, this statistic tends to be larger in distribution than when the watermark is absent. One can conclude that a watermark is detected if this statistic is above a certain threshold. Despite being intuitive, however, this statistic is ad-hoc, and one cannot rule out the possibility of a better detection rule.
To better understand practical watermarking schemes, we turn to real-world LLMs where the size of the token vocabulary is very large. As a recap, at the core of an unbiased watermark is a sampling method for multinomial distributions. Perhaps the two most common sampling methods are the Gumbel-max trick [42, 32] and the inverse transform [19]. Interestingly, these two sampling methods correspond precisely to two important and representative watermarks [38, 1].
Gumbel-max watermark [1]
Let $\zeta=(U_{w})_{{w}\in{{\mathcal{W}}}}$ consist of $|{{\mathcal{W}}}|$ i.i.d. copies of $U(0,1)$. A version of the Gumbel-max trick [26] states that $\arg\max\limits_{{w}\in{{\mathcal{W}}}}\frac{\log U_{w}}{P_{w}}$ follows the NTP distribution $\bm{P}\equiv(P_{{w}})_{{w}\in{{\mathcal{W}}}}$. [^5] It is noteworthy that the Gumbel-max trick has a broad range of applications, including partition function estimation [28], ranking under the random utility model [57], and computational sampling in machine learning [48, 42, 32]. Recognizing this fact, Scott Aaronson proposed the following decoder [1]:
@@
Inverse Transform Watermark [38]
It is well-known that any univariate distribution can be sampled by applying the inverse cumulative distribution function (CDF) to $U(0,1)$. Given an NTP distribution $\bm{P}$ and $\pi$ that maps all tokens in ${{\mathcal{W}}}$ to a permutation of $1,2,\ldots,|{{\mathcal{W}}}|$, consider the multinomial distribution with probability mass $P_{\pi^{-1}(i)}$ at $i$ for $1\leq i\leq|{{\mathcal{W}}}|$. The CDF of this distribution takes the form $F(x;\pi)=\sum_{{w}^{\prime}\in{{\mathcal{W}}}}P_{{w}^{\prime}}\cdot{\mathbf{1}}_{\{\pi({w}^{\prime})\leq x\}}.$ Taking as input $U\sim U(0,1)$, the generalized inverse of this CDF is defined as $F^{-1}(U;\pi)=\min\{i:\sum_{{w}^{\prime}\in{{\mathcal{W}}}}P_{{w}^{\prime}}\cdot{\mathbf{1}}_{\{\pi({{w}^{\prime}})\leq i\}}\geq U\},$ which by construction follows the multinomial distribution $\bm{P}$ after applying the permutation $\pi$. Making use of this fact, [38] proposed the inverse transform watermark with the following decoder:
@@
where the pseudorandom variable $\zeta:=(\pi,U)$ and the permutation $\pi$ is uniformly at random. By definition, this watermark is unbiased. To detect the watermark, [38] proposed detection rules that, roughly speaking, examine the absolute difference between $U_{t}$ and the rank of the generated token, $\pi_{t}({w}_{t})$, and sum the differences across the token sequence (see details in Section 4). When the text is watermarked, the difference turns out to be small because the involved two random variables are highly dependent. Similar to the Gumbel-max watermark, the detection of this watermark could potentially benefit from a more principled derivation to enhance its power.
@@ 1.1 This Paper
From a statistical viewpoint, a watermark scheme’s performance at the detection phase is evaluated by two competing criteria. The first is how likely it would mistakenly detect human-written text as LLM-generated text, and the second is the probability that it would mistakenly classify LLM-generated text as its human-written counterpart. This perspective relates watermarks to hypothesis testing, which formally calls the two aforementioned errors Type I error and Type II error, respectively. In this regard, the effectiveness of the Gumbel-max watermark and inverse transform watermark, even including the baby watermark, is not clear yet, though [22, 49] compared their efficiency through empirical experiments. For example, even though they come with detection rules that seem intuitive, it is not clear if they are statistically optimal in the sense of having the optimal trade-off between Type I and Type II errors. If they are not optimal, it is of interest to find a better detection rule.
@@ 1.2 Related Work
The most closely related work to ours is the Gumbel-max watermark [1] and the inverse transform watermark [38]. In [78, 22], the authors introduced unbiased watermarks that are robust to probability perturbations or multi-bit processing. Other unbiased watermarks in this fast-growing line of research include [15, 70, 30]. In addition, a popular example is the red-green list watermark, which splits the vocabulary into red-green lists based on hash values of previous n-grams and slightly increases the probability of green tokens embedding the watermark [35, 36, 77, 41, 12]. In the detection phase, a high frequency of green tokens suggests the text is LLM-generated. This type of watermark is biased since the NTP distributions have been altered, thereby leading to performance degradation of the LLM.
In contrast, there is much less work on the theoretical front. A notable exception is the work of [31], which approached watermark detection from the perspective of composite independence testing. Their framework requires model providers to offer random rejection regions. In contrast, our approach employs a pivotal statistic to detect distributional shifts between human-written and LLM-generated text, allowing verifiers to choose any rejection region. While both approaches aim to minimize a form of worst-case Type II error, the optimal detection rule proposed by [31] is hindered by a non-vanishing Type II error due to the consideration of many worst-case scenarios, and it is computationally inefficient because of the exponentially large rejection regions required. [^6] See Theorem 3.10 of [31] and the subsequent discussion. Conversely, our optimal rule generally achieves a vanishing Type II error by addressing fewer worst-case scenarios, benefiting from the well-regularized null behavior of pivotal statistics, and is computationally efficient due to its sum-based structure.
@@ 2.1 Working Hypotheses
Working Hypothesis 2.1 is grounded purely in cryptographic considerations. In cryptography, there are well-established approaches to efficiently constructing and computing the pseudorandom number as a function of text and the secret key [8, 47, 59, 55, 33]. The pseudorandom number is very sensitive to the key, making it computationally indistinguishable from its truly random counterpart without knowledge of the key. [^7] The phrase “computationally indistinguishable” means that no polynomial-time algorithm can distinguish the pseudorandom number from its truly random counterpart without knowledge of the key. Specifically, although $\zeta_{t}$ is completely determined by the prior text ${w}_{1:(t-1)}$ (and the secret key), a run of the hash function ${\mathcal{A}}$ could effectively introduce fresh randomness, thereby making $\zeta_{t}$ statistically independent of ${w}_{1:(t-1)}$. Note that, during detection, the communication cost is merely to pass the secret key to the verifier as the prior tokens are publicly available.
@@
The two working hypotheses have been adopted by the literature, albeit not as explicit as our treatment. For example, [70, 78] assumed conditions on the pseudorandom hash function so that the first working hypothesis is valid. [^8] In particular, they assumed that ${\mathcal{A}}({w}_{1:n},{\mathtt{Key}})$ is i.i.d. for any ${w}_{1:n}$. Other works directly impose distributional assumptions on some summary statistics to effectively satisfy the two working hypotheses [35, 36, 77, 22]. As a departure from these works, [38] considered a hash function of the form $\zeta_{t}={\mathcal{A}}(t,{\mathtt{Key}})$. Consequently, the independence between the pseudorandom number and prior tokens follows by construction. However, it is worthwhile mentioning that this form of the hash function hampers computational efficiency in detection due to the necessity of iterating through the entire text sequence multiple times.
@@ 2.2 Pivotal Statistics
That is, we reject that the text is written by a human if $T_{h}$ is above $\gamma_{n,\alpha}$. The threshold $\gamma_{n,\alpha}$ is chosen to ensure significance level at $\alpha$: ${\mathbb{P}}_{H_{0}}(T_{h}(Y_{1:n})=1)=\alpha$. For certain score functions discussed in this work, the distribution of the sum of $Y_{t}$’s under the null hypothesis admits a closed-form expression, allowing for analytical evaluation of $\gamma_{n,\alpha}$. In general, as the null distribution of $h\circ Y$ is known, an estimator of $\gamma_{n,\alpha}$ when the text is sufficiently long is
@@ 2.3 Class-Dependent Efficiency
Once the Type I error is controlled, we seek to evaluate the Type II error and use it as a measure to ascertain which choice of score function is more desired than others. If the NTP distributions $\bm{P}_{t}$ were known and remained the same with respect to varying $t$, the optimal score function would simply be given by the log-likelihood ratio, according to the Neyman–Pearson lemma. However, this is not the case. We manage this challenge by assuming that the NTP distributions belong to a distribution class, denoted as ${\mathcal{P}}$. The flexibility of using a distribution class lies in that one can choose a small class when much is known about the distributional properties of the LLM, and choose a large class if little is known.
@@
$R_{{\mathcal{P}}}(h)$ is the ${\mathcal{P}}$-dependent efficiency rate of $h$, or simply the ${\mathcal{P}}$-efficiency rate.
@@
Under a certain regularity condition, (12) is tight in the sense that there exists $\bm{P}^{\star}$ in ${\mathcal{P}}$ such that, if $\bm{P}_{t}=\bm{P}^{\star}$ for all $t$, then ${\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)\geq\mathrm{e}^{-(R_{{\mathcal{P}}}(h)+\varepsilon)n}$ for any positive $\varepsilon$ and sufficiently large $n$. The regularity condition is satisfied by both the Gumbel-max and inverse transform watermarks and is detailed in the remark following the proof of Theorem 2.1 in Appendix A.
@@
This efficiency rate, obtained from large deviation theory, does not depend on the specific value of $\alpha$. The same is true for Bahadur efficiency [6], which quantifies the rate at which the significance level ($p$-value) of a test statistic approaches zero with increasing sample size. It emphasizes the control of Type I errors for simple hypothesis testing with i.i.d. data. In contrast, our metric $R_{{\mathcal{P}}}(h)$ centers on quantifying the least decline of Type II errors to zero when the underlying NTP distribution $\bm{P}$ is picked up from the distribution class ${\mathcal{P}}$. In short, it focuses on composite hypothesis testing where the observed data is not identically distributed.
Optimality via minimax optimization
@@
$\sup_{h}R_{{\mathcal{P}}}(h)$ $=-\inf_{h,\theta\geq 0}\sup_{\bm{P}\in{\mathcal{P}}}\left(\operatorname{\mathbb{E}}_{0}\theta h(Y)+\log\operatorname{\mathbb{E}}_{1,\bm{P}}\mathrm{e}^{-\theta h(Y)}\right).$
@@
Choice of distribution classes
@@
Assume that at least $\gamma$-fraction of $\bm{P}_{1},\ldots,\bm{P}_{n}$ is in ${\mathcal{P}}_{1}$ with the rest being in ${\mathcal{P}}_{2}$, [^12] We do not need to know which are in the $\gamma$ fraction. where $10<\gamma<$ 1 and ${\mathcal{P}}_{1}\subset{\mathcal{P}}_{2}$. Then, for any $h$, we have
$\limsup\limits_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)^{1/n}\leq\mathrm{e}^{-\{\gamma\cdot R_{{\mathcal{P}}_{1}}(h)+(1-\gamma)\cdot R_{{\mathcal{P}}_{2}}(h)\}}.$
When there is no prior at all about the LLMs, the associated class is the $(|{{\mathcal{W}}}|-1)$-dimensional simplex, which contains all possible NTP distributions and we denote by ${\mathrm{Simp}}({{\mathcal{W}}})$. This distribution class includes singular distributions—a distribution $\bm{P}$ such that $P_{{w}}=1$ for some token ${w}$—for which the pivotal statistic has the same distribution under the null and alternative in (9). Formally, this gives $R_{{\mathrm{Simp}}({{\mathcal{W}}})}(h)=0$, that is, the class-dependent efficiency of ${\mathrm{Simp}}({{\mathcal{W}}})$ is zero. Taking ${\mathcal{P}}_{1}={\mathcal{P}}\subset{\mathcal{P}}_{2}={\mathrm{Simp}}({{\mathcal{W}}})$, Proposition 2.1 shows $\lim\sup\limits_{n\to\infty}{\mathbb{P}}_{H_{1}}(T_{h}(Y_{1:n})=0)^{1/n}\leq\mathrm{e}^{-\gamma R_{{\mathcal{P}}}(h)}.$ This result implies that when only a fraction of NTP distributions is known to some extent, the Type II error is still governed by $R_{{\mathcal{P}}}(h)$ up to a multiplicative constant $\gamma$. The comparison between choices of the score functions $h$ can still be reduced to comparing the value of $R_{{\mathcal{P}}}(h)$.
@@
the $\Delta$-regular distribution class. Accordingly, an NTP distribution $\bm{P}$ is called $\Delta$-regular if $\bm{P}\in{\mathcal{P}}_{\Delta}$. This excludes singular distributions. Interestingly, $\Delta$-regularity is closely related to the Shannon entropy. For any $\Delta$-regular NTP distribution $\bm{P}$, its Shannon entropy satisfies ${\mathrm{Ent}}(\bm{P})=\sum P_{{w}}\log\frac{1}{P_{w}}\geq\sum P_{{w}}(1-P_{w})\geq\sum P_{{w}}\cdot\Delta=\Delta$, where the first inequality follows because $\log\frac{1}{1-x}\geq x$ for $x<1$. This shows that $\Delta$-regularity imposes a lower bound on how much information the NTP distribution offers. In practice, most NTP distributions are $\Delta$-regular for a proper value of $\Delta$. See Figure 1 for an empirical investigation where 88.65% of the NTP distributions are 0.0001-regular, 81.36% are 0.001-regular, and 70.28% are 0.01-regular.